OpenStudy (anonymous):

Simplify the final answer: (4 + sqrt(7)) * (4 - sqrt(7))

OpenStudy (saifoo.khan):

FOIL.

OpenStudy (anonymous):

9.

OpenStudy (saifoo.khan):

\[(4 + \sqrt{7}) * (4 - \sqrt{7}) \]

OpenStudy (anonymous):

16-7=9.

myininaya (myininaya):

\[(a-b)(a+b)=a^2-b^2\]

myininaya (myininaya):

\[(4)^2-(\sqrt{7})^2=16-7=9 \]

OpenStudy (anonymous):

as we know that \[(a-b)(a+b)=a^2-b^2\] hence \[(4+sqrt7)(4-sqrt7)\] \[((4)^2-(sqrt7)^2\] \[16-7\] since 4^2=16 and sqrt7^2=7 \[9\] got it

OpenStudy (anonymous):

did you really say FOIL?

OpenStudy (anonymous):

thanks kushashwa23 - makes complete sense!